We want to pump water from a well 20 m deep at the rate of 500 L/min. Assuming 60 % pump efficiency, how powerful should the pump motor be?



1
Expert's answer
2021-02-15T00:57:05-0500

The pump motor power can be found as follows:


"P=\dfrac{Q\rho gh}{\eta_t},"

here, "Q" is the flow rate of the pump, "\rho" is the density of the water, "g" is acceleration due to gravity, "h" is the head of the pump, "\eta_t" is the pump efficiency.

Then, we get:


"P=\dfrac{500\ \dfrac{L}{min}\cdot\dfrac{1\ min}{60\ s}\cdot\dfrac{10^{-3}\ m^3}{1\ L}\cdot1000\ \dfrac{kg}{m^3}\cdot9.8\ \dfrac{m}{s^2}\cdot20\ m}{0.6}=2722\ W."